OpenStudy (anonymous):

I'm not understanding linear system by substitution. Can I please have an explanation on how to solve this problem? 5y = 3x + 15 6x = 10y - 30

OpenStudy (anonymous):

If you multiply the first expression by 2 you will get- 6x + 10y = 30 6x + 10y = -30 Since the equations intercepts are different, and they have same slopes... they don't intersect each other. Therefor there is no solution.

OpenStudy (uri):

5y = 3x + 15 6x = 10y - 30 Re arrange the equations first -3x+5y=15(multiply by 2) 6x-10y=-30 -6x+10y=30 6x-10y=-30 ========== -6x+6x will be cancelled. +10y-10y will be cancelled. +30-30 will be cancelled. So no solution.

OpenStudy (anonymous):

Thank you so much for the step by step. Preciate it. @uri

OpenStudy (uri):

Yw :)